# HyperNet Fields: Efficiently Training Hypernetworks without Ground Truth by Learning Weight Trajectories

### v1

@@ 1 Introduction
Various generative modeling tasks [36] require adaptation, thus data-dependent network weights, or in other words, model parameters being adjusted according to various circumstances (conditions and user input). These include personalized text-to-image generation where the weights of the underlying personalization model such as DreamBooth [36] are fine-tuned on the concept of interest and used to generate target images conditioned on input prompts. Another example is modeling 3D, where neural networks are used to model shapes implicitly with signed distance fields [27] or occupancy [25]. In both cases, a dedicated set of neural network weights is required for each personalization outcome or 3D shape. Thus, the application of these methods is highly limited by the training time required to obtain these weights.
@@ 2 Related Work
Historically, hypernetworks have been trained by backpropagating gradients through the task-specific network to the hypernetwork itself, as demonstrated by early approaches in hypernetworks [14] and subsequent explorations in continual learning [26] and neural architecture search [46]. These methods have shown that allowing gradient flow through both networks enables the hypernetwork to adaptively modify parameters for diverse tasks. However, more recent works have introduced approaches that involve precomputing weights for the task-specific network and directly supervising the hypernetwork to match them, eliminating the need for backpropagation through the task-specific network and improving training stability by alleviating issues with gradient flow. This precomputed weight supervision has been shown to enhance gradient stability and training performance, as highlighted in frameworks like HyperDiffusion [10] and HyperDreamBooth [37], however this limits the size of dataset this can be applied to as the precompute requirement is proportional to the number of samples in the dataset. Our method combines the best of both approaches by removing the need for precomputed weights while also explicitly providing weight based supervision in the form of gradients along the convergence path, which enables stable training on large datasets.
@@
In generative modeling, finding smooth and reliable paths through latent space has become central to producing coherent samples. Examples include diffusion models [40, 15] which iteratively trace a path from noise to data through successive steps, effectively navigating high-dimensional space by reversing a stochastic process. Consistency models [43] refine this by learning non-Markovian trajectories that reach target data distributions in fewer steps by introducing a formulation that allows them to single-step jump to any timestep. Flow-matching models [21] take a similar path-based approach by learning a vector field that guides samples along pre-defined flows, ensuring a direct and smooth transition through the latent space. Similarly to these methods, our method can be thought of as finding a path through this high dimensional space, specifically in the form of a single-step estimator from initialization to any desired state of convergence, whose path travels strictly along the direction of steepest gradient descent.
@@
Several recent approaches in generative modeling and physics-informed learning employ supervision exclusively through gradients, rather than direct supervision between the input-output pairs, to achieve denoising and reconstruction. Score-based generative models [41, 44, 42] use score matching to learn the gradient (or ”score”) of the log density of noisy data distributions, bypassing the need for direct output supervision. Denoising Score Matching (DSM) [44] trains the model to predict this gradient, guiding it to approximate data distributions through iterative denoising and demonstrating improved stability in noise removal. Similarly, Physics-Informed Neural Networks (PINNs) [34, 32, 33] leverage known derivatives from partial differential equations (PDEs) in physical simulations, such as heat diffusion or fluid flow, to inform the network solely through spatial or temporal derivatives. This derivative-based guidance encourages convergence as the model adapts to match observed physical behavior, requiring no explicit target outputs. Energy-Based Models (EBMs) [12, 20, 9] also align with this gradient-driven approach by learning energy landscapes that adjust gradients to represent target distributions, allowing for smooth gradient matching through learned energy fields. Inspired by these methods, our approach supervises only the gradient direction in task-specific weight space, effectively learning a convergence pathway without requiring explicit final state supervision. By supervising gradients within this pathway, our method achieves robust training stability while generalizing over diverse tasks without per-sample ground truth weights.
@@ 3.2 Gradient-based Supervision
Forcing 0 offset at t=0 by construction
@@ 4 Experiments
We consider two diverse applications—image personalization and 3D shape reconstruction as the target applications. While we focus on these two, we emphasize that our hypernet fields framework does not make any assumption on the task-specific network and applies to any task. Moreover, the definition of the loss, $L_{task}(\theta_{t},c)$ is not restricted to a particular domain. Thus, our method can be applied to these problems by just modifying the task-specific optimization model.
@@ 4.1 Application to Personalized Image Generation
Figure 4: Qualitative examples (AHFQ) – Qualitative results of personalized animal image generation from the AFHQ dataset [6] demonstrate the adaptability of our hypernetwork field. Compared to DreamBooth [36] and Textual Inversion [11], our method effectively captures specific animal’s characteristics while enabling rapid adaptation. Further examples are available in the supplementary material.
Task network – DreamBooth
@@
Note that, unlike existing works that require per-instance fitting, with our method, it is possible to introduce on-the-fly data augmentation. We thus incorporate color, saturation, contrast, and hue augmentations to perturb the appearance of the conditioning signal $\mathbf{x}$, as well as introduce horizontal flips.
Evaluation metrics
@@ Results.
We provide qualitative highlights in Figs. 3 and 4, and the quantitative summaries in Tables 1 and 2. Our results for image personalization in Figs. 3 and 4 is on par with the target Dreambooth model whose weights we estimate with our HyperNet Fields approach. That is, our approach generates personalized images that are representative of the given input prompt and the conditional image. Again, note that we achieve these personalization results in a single step by querying the hypernetwork once at timestep $T$. Interestingly, our method demonstrates remarkable performance for preserving the original image context, as measured by DINO and CLIP-I scores (Tab. 1). Further, in the case of the AFHQ dataset (Tab. 2), we outperform all DreamBooth variants, demonstrating that not only do we remove the compute-heavy optimization requirement of DreamBooth [36], but also by amortizing the behavior of the learned weights with respect to the input conditioning image, we successfully replicate the dynamics of the task-specific dreambooth network yielding improved results. We emphasize once more that while results are comparable to each other, our method is significantly faster than the baselines.
@@ 5 Conclusion
Our method enables stable hypernetwork training on very large datasets. Because of this, the potential of our method for diverse tasks remains underexplored. This may include tasks such as training a hypernetwork to estimate the weights of NeRFs, or even personalized adaptors for large language models. These would be interesting avenues for future work.

### v2

@@ 1 Introduction
Various generative modeling tasks [40] require adaptation, thus data-dependent network weights, or in other words, model parameters being adjusted according to various circumstances (conditions and user input). These include personalized text-to-image generation where the weights of the underlying personalization model, such as DreamBooth [40], are fine-tuned on the concept of interest and used to generate target images conditioned on input prompts. Another example is 3D modeling, where neural networks are used to model shapes implicitly with signed distance fields [30] or occupancy [29]. In both cases, a dedicated set of neural network weights is required for each personalization outcome or 3D shape. Thus, the applicability of these methods is highly limited by the training time required to obtain these weights.
@@ 2 Related Work
Early work on training hypernetworks backpropagates gradients through the task-specific network to the hypernetwork itself, as demonstrated by approaches [16, 32] and subsequent explorations in continual learning [49] and neural architecture search [51]. These methods have shown that allowing gradient flow through both networks enables the hypernetwork to modify parameters for diverse tasks adaptively. However, more recent works have introduced approaches that involve precomputing weights for the task-specific network and directly supervise the hypernetwork to match them, eliminating the need for backpropagation through the task-specific network and improving training stability by alleviating issues with gradient flow. This supervision with precomputed weights has been shown to enhance gradient stability and training performance, as highlighted in frameworks like HyperDiffusion [11], HyperFields [2], and HyperDreamBooth [41]. However, this limits the size of the dataset to which this can be applied, as the precompute requirement is proportional to the number of samples in the dataset. Our method combines the best of both approaches by removing the need for precomputed weights while explicitly providing weight-based supervision in the form of gradients along the convergence path, enabling stable training on large datasets.
@@
In generative modeling, finding smooth and reliable paths through latent space has become central to producing coherent samples. Examples include diffusion models [44, 17], which iteratively trace a path from noise to data through successive steps, effectively navigating high-dimensional space by reversing a stochastic process. Consistency models [47] refine this by learning non-Markovian trajectories that reach target data distributions more efficiently, introducing a formulation allowing single-step transitions to any timestep; subsequent works [23, 12] further improve this by supervising the entire denoising trajectory with guidance from a teacher model. Flow-matching models [25] take a similar path-based approach by learning a vector field that guides samples along pre-defined flows, ensuring a direct and smooth transition through the latent space. Similarly to these methods, our method can be thought of as finding a path through this high-dimensional space, specifically in the form of a single-step estimator from initialization to any desired state of convergence, whose path travels strictly along the direction of the steepest gradient descent.
@@
Several recent approaches in generative modeling and physics-informed learning employ supervision exclusively through gradients rather than direct supervision between input-output pairs to achieve denoising and reconstruction. Score-based generative models [45, 48, 46] use score matching to learn the gradient (or ”score”) of the log density of noisy data distributions, bypassing the need for direct output supervision. Denoising Score Matching (DSM) [48] trains the model to predict this gradient, guiding it to approximate data distributions through iterative denoising and demonstrating improved stability in noise removal. Similarly, Physics-Informed Neural Networks (PINNs) [38, 36, 37] leverage known derivatives from partial differential equations (PDEs) in physical simulations, such as heat diffusion or fluid flow, to inform the network solely through spatial or temporal derivatives. This derivative-based guidance encourages convergence as the model adapts to match observed physical behaviour, requiring no explicit target outputs. Energy-Based Models (EBMs) [14, 24, 10] also align with this gradient-driven approach by learning energy landscapes that adjust gradients to represent target distributions, allowing for smooth gradient matching through learned energy fields. Inspired by these methods, our approach supervises only the gradient direction in task-specific weight space, effectively learning a convergence pathway without requiring explicit final state supervision. By supervising gradients within this pathway, our method achieves robust training stability while generalizing over diverse tasks without per-sample ground truth weights.
@@ 3.2 Gradient-based Supervision
Forcing 0 offset at t=0 by construction.
@@ 4 Experiments
We consider two diverse applications – image personalization and 3D shape reconstruction as target applications. While we focus on these two, we emphasize that our hypernet fields framework does not make any assumption on the task-specific network and applies to any task. Moreover, the definition of the loss, $L_{task}(\theta_{t},c)$ is not restricted to a particular domain. Thus, our method can be applied to these problems by modifying the task-specific optimization model.
@@ 4.1 Application to Personalized Image Generation
Figure 4: Qualitative examples (AHFQ) – Qualitative results of personalized animal image generation from the AFHQ dataset [6] demonstrate the adaptability of our hypernetwork field. Compared to DreamBooth [40] and Textual Inversion [13], our method effectively captures the specific animal’s characteristics while enabling rapid adaptation. Further examples are available in the supplementary material.
Task network – DreamBooth.
@@
Note that, unlike existing works that require per-instance fitting, with our method, it is possible to introduce on-the-fly data augmentation. We thus incorporate color, saturation, contrast, and hue augmentations to perturb the appearance of the conditioning signal $\mathbf{x}$, including horizontal flips.
Evaluation metrics.
@@ Results.
We provide qualitative highlights in Figs. 3 and 4, and the quantitative summaries in Tables 1 and 2. Our results for image personalization in Figs. 3 and 4 is on par with the target Dreambooth model whose weights we estimate with our HyperNet Fields approach. That is, our approach generates personalized images that are representative of the given input prompt and the conditional image. Again, note that we achieve these personalization results in a single step by querying the hypernetwork once at timestep $T$. Interestingly, our method demonstrates remarkable performance for preserving the original image context, as measured by DINO and CLIP-I scores (Tab. 1). Further, in the case of the AFHQ dataset (Tab. 2), we outperform all DreamBooth variants, demonstrating that not only do we remove the compute-heavy optimization requirement of DreamBooth [40], but also by amortizing the behavior of the learned weights with respect to the input conditioning image, we successfully replicate the dynamics of the task-specific dreambooth network yielding improved results. We emphasize that while results are comparable to each other, our method is significantly faster than the baselines.
@@ 5 Conclusion
Our method enables stable hypernetwork training on large datasets. Because of this, the potential of our method for diverse tasks remains underexplored. This may include tasks such as training a hypernetwork to estimate the weights of NeRFs or even personalized adapters for large language models. However, our approach requires as many output neurons as there are weights in the target network, resulting in suboptimal memory usage. Recent work by Dravid et al. [9] has shown that personalized DreamBooth adapters can be represented using only 1k parameters. Integrating our method with this compact representation approach could significantly enhance parameter efficiency while maintaining our proposed benefits of efficient hypernetwork training. These would be interesting avenues for future work.